# BayTTA: Uncertainty-aware medical image classification with optimized test-time augmentation using Bayesian model averaging

### v1

@@ 4.3.2 Results for gene editing datasets
Fig. 5: Comparison of the TTA and BayTTA method performance on the skin cancer dataset in terms of accuracy and standard deviation, while considering pre-trained baseline models with rotate, zoom, and shift augmentations.
Fig. 6: Comparison of the TTA and BayTTA method performance on the breast cancer dataset in terms of accuracy and standard deviation, while considering the pre-trained baseline models with rotate, zoom, and shift augmentations.
Fig. 7: Comparison of the TTA and BayTTA method performance on the chest X-ray dataset in terms of accuracy and standard deviation, while considering the pre-trained baseline models with rotate, zoom, and shift augmentations.
In addition to the comparison with pre-trained baseline models, we conducted a comparative analysis between BayTTA and TTA used in combination with state-of-the-art DL classification models whose input was adapted to gene editing datasets. The results obtained in this experiment are presented in Tables 10 and 10. To conduct our assessment, we used original images with ten randomly created samples (generated for each original image during the testing phase using rotations and random erasing). The results reported in Table 10 suggest that the proposed BayTTA computational framework used in combination with the COVID-19 model provided the highest accuracy and the lowest STD values for both CRISPOR and GUIDE-seq datasets. Similar model performances can be observed in Table 10, where the obtained precision, recall, and F1-score metric values are reported - the COVID-19+BayTTA model provided the best results overall.
4.4 Assessing the impact of different data augmentations and of increasing the number of samples
@@
Fig. 8: Comparison of the TTA and BayTTA method performance in terms of accuracy with different numbers of samples, while considering the best baseline model from Tables 4, 4, and 4 for each medical image dataset.
4.5 Evaluation of different data augmentations
@@
Our results are graphically represented in Figures 5, 6, and 7, corresponding to the skin cancer, breast cancer, and chest X-ray datasets, respectively. These graphs demonstrate how the accuracy and STD of the model alter as we vary data augmentation types, highlighting which augmentations offer higher accuracy and boost the robustness of CNN models.
@@
4.6 Evaluation of increasing the number of samples
We analyzed the impact of increasing the number of augmented samples on the models accuracy during the testing phase. As indicated in Tables 4, 4 and 4, the combination of BMA and TTA contributes to a higher accuracy of DenseNet201 on the skin cancer dataset, of ResNet152V2 on the breast cancer dataset, and of VGG-16 on the chest X-ray dataset, compared to baseline models and TTA. Thus, we examined the effect of increasing the number of augmented samples in these three specific cases.
Figure 8 demonstrates that the accuracy of both TTA and BayTTA increases as the number of samples in the test data with diverse transformations grows. Hence, training the model to recognize patterns in their diverse forms generally enhances its ability to make accurate predictions on unseen examples. Obviously, the increase in the amount of augmented data results in a higher accuracy at the expense of a higher computational cost.
5 Conclusion

### v2

@@ IV-C2 Results for gene editing datasets
Figure 5: Comparison of the TTA and BayTTA method performance on the skin cancer dataset in terms of accuracy and standard deviation, while considering pre-trained baseline models with rotate, zoom, and shift augmentations.
Figure 6: Comparison of the TTA and BayTTA method performance on the breast cancer dataset in terms of accuracy and standard deviation, while considering the pre-trained baseline models with rotate, zoom, and shift augmentations.
Figure 7: Comparison of the TTA and BayTTA method performance on the chest X-ray dataset in terms of accuracy and standard deviation, while considering the pre-trained baseline models with rotate, zoom, and shift augmentations.
In addition to the comparison with pre-trained baseline models, we conducted a comparative analysis between BayTTA and TTA used in combination with state-of-the-art DL classification models whose input was adapted to gene editing datasets. The results obtained in this experiment are presented in Tables X and X. To conduct our assessment, we used original images with ten randomly created samples (generated for each original image during the testing phase using rotations and random erasing). The results reported in Table X suggest that the proposed BayTTA computational framework used in combination with the COVID-19 model provided the highest accuracy and the lowest STD values for both CRISPOR and GUIDE-seq datasets. Similar model performances can be observed in Table X, where the obtained precision, recall, and F1-score metric values are reported - the COVID-19+BayTTA model provided the best results overall.
IV-D Assessing the impact of different data augmentations and of increasing the number of samples
@@
IV-E Evaluation of different data augmentations
@@
Our results are graphically represented in Figs. 5, 6, and 7, corresponding to the skin cancer, breast cancer, and chest X-ray datasets, respectively. These graphs demonstrate how the accuracy and STD of the model alter as we vary data augmentation types, highlighting which augmentations offer higher accuracy and boost the robustness of CNN models.
@@
Figure 8: Comparison of the TTA and BayTTA method performance in terms of accuracy with different numbers of samples, while considering the best baseline model from Tables IV, IV, and IV for each medical image dataset.
IV-F Evaluation of increasing the number of samples
We analyzed the impact of increasing the number of augmented samples on the models accuracy during the testing phase. As indicated in Tables IV, IV and IV, the combination of BMA and TTA contributes to a higher accuracy of DenseNet201 on the skin cancer dataset, of ResNet152V2 on the breast cancer dataset, and of VGG-16 on the chest X-ray dataset, compared to baseline models and TTA. Thus, we examined the effect of increasing the number of augmented samples in these three specific cases.
Fig. 8 demonstrates that the accuracy of both TTA and BayTTA increases as the number of samples in the test data with diverse transformations grows. Hence, training the model to recognize patterns in their diverse forms generally enhances its ability to make accurate predictions on unseen examples. Obviously, the increase in the amount of augmented data results in a higher accuracy at the expense of a higher computational cost.
V Conclusion